The Dixie Chicken currently sells 3,000 burger platters per month for $7, and the own-price elasticity for this platter has been estimated to be -1.3. If Dixie Chicken raises prices by 70 cents, how many platters will be sold?
Answer #1

E=-1.3

% increase in price = .7/7 = 10%

So,

-1.3 = % change in quantity demanded / % increase in price

% change in quantity demanded = -1.3*10% = 13%

So, new number of platters, sold will decrease by 13%.

No. of platters to be sold, now = 3000*(1-13%)

No. of platters to be sold, now = 2610